# Information Discovery in e-Commerce

### v1

@@ 6.2.2 Embedding-based methods
Embedding-based methods are another type of candidate retrieval methods. As shown in Fig. 6.3, this kind of method first maps users and items into a common embedding space, and then leverages approximate the K-Nearest Neighbor (KNN) algorithm [andoni2006near] to search for candidate items with the smallest embedding distance to the target user. The key of these methods lies in learning high-quality embeddings for each user and item. In what follows, we introduce techniques for learning embeddings.
Matrix factorization. Matrix factorization (MF) is a classic embedding-based method. The basic assumption behind MF is that the user-item interaction matrix has a low-rank structure. As such, MF delineates each user and item as an embedding vector and then predicts the preference between each user-item pair as the inner product of their embedding vectors. Let $p_{u}\in\mathcal{R}^{k}$, $q_{i}\in\mathcal{R}^{k}$ denote the embedding vector of user $u$ and item $i$, respectively. MF makes a prediction for the user-item pair $(u,i)$ as follows:
@@
Compared with heuristic-based methods, MF is a more generic method that can adaptively learn user preferences from their history behaviors and require no manually crafted heuristic design. The use of MF has brought a revolution, drawing research attention from previous heuristic-based methods towards recent embedding-based methods.
Information-enhanced embedding models. Despite the prevalence of MF, it is still insufficient to yield accurate embeddings. The reason is that MF directly projects user/item IDs to an embedding space, making MF reliant on the behavioral signal from the objective function. Hence, MF models do not perform well for inactive users or items with very few interactions. To deal with this problem, several methods have been proposed to enrich the representation with supplementary information. We can divide them into three groups: (1) neighborhood-enriched embedding methods (2) feature-enriched embedding methods and (3) graph-enriched embedding methods. We discuss each of these groups in detail as follows:
(1) Neighborhood-enriched embedding. Beyond user ID, bell2007lessons, koren2008svd++, kabbur2013fism enrich a user’s representation with their rating history. This kind of method generates user-item preference scores as follows:
@@
(2) Feature-enriched embedding. Another way to improve MF is to leverage rich user and item features, e.g., a user’s age, gender, education, revenue, product tags, category, and price. These features are valuable in enriching the representations of users and items, which further boosts the recommendation performance. Fig. 6.4 summarizes the architecture of feature-enriched candidate retrieval methods as a two-tower structure [yi2019youtube-2tower, fan2019mobius, xu2016tag-dssm]. The left tower is the user tower that translates a user’s features into the user’s embedding representation; the right tower is the item tower generating the item representation. Let $x_{u}$ denote the features of user $u$, while $x_{i}$ denotes the features of item $i$. The function of the two towers can be depicted as follows:
@@
where $f_{U}(\cdot)$ and $f_{I}(\cdot)$ denote the translation function with regard to the implementation of the two towers. Linear models and neural networks (e.g., MLP [huang2013dssm], LSTM [song2016TDDSSM], CNN [shen2014CDSSM] and auto-encoders [CDAE, liang2018variational]) can be used as the translation function. Given user and item embeddings, the two-tower model makes a prediction for each user-item pair. Cosine similarity and inner product are usually adopted to measure the embedding distance, so we have:
@@
(3) Graph-enriched embedding. A drawback of the aforementioned methods is that they regard each user or item as an “island” and fail to explicitly encode their relations into representations. In fact, these relations, e.g., item-item occurrences or user-item interactions, are important in revealing correlations between users and items, which provide valuable signals for recommendation. To address this problem, several studies have established specific graphs between users and items, and then conduct graph representation learning to generate user and item embeddings. One representative method is EGES [EGES]. EGES constructs an item-item graph based on user behavior sequences, where two items are connected if they consecutively occur in one sequence. EGES then applies DeepWalk [perozzi2014deepwalk] on the item graph to generate item representations. Similarly, other related methods perform graph convolutional networks to enrich the representation learning of both users and items [wang2019NGCF, pingsage, LightGCN]. The knowledge graph with informative relations between items has been exploited to learn better embeddings [wang2019kgat, wang2020ckan].
@@ 6.2.3 Session-based recommendation
Early studies on sequential recommendation methods often use Markov chains by assuming that the user’s next action only depends on the previous one. As a representative method, the factorizing personalized Markov chains model (FPMC) extends MF by modeling the effects of sequential-consecutive actions [FPMC]. Since the previous action is indeed a critical factor affecting the user’s next decision, FPMC achieves a significant gain over MF-based models. Following FPMC, he2016fusing leverage a higher-order Markov chain to capture the sequential dependence among non-consecutive behavior. it is hard to use Markov chain-based methods for capturing complicated and long-term dependencies in sequential data, and this limits their performance in e-commerce recommendation.
More recently, deep neural networks have been leveraged in sequential recommendation due to its powerful expressive ability on capturing behavior dependences (discussed in Chapter 4). Generally speaking, neural networks for sequential recommendation can be divided into three types: (i) RNN-based methods that model sequential dependence with RNN (or improved versions such as LSTM and GRU) to capture both long-term and short-term dependencies [quadrana2017personalizing, jannach2017recurrent, hidasi2016parallel]; (ii) CNN-based methods that concatenates the embedding of previous item in a sequence as to a matrix [tang2018personalized]; and (iii) attention-based methods that introduce attention mechanism into sequential recommendation by considering various types of user behaviors with different influences [kang2018self, sun2019bert4rec].
@@ 6.2.4 Next-basket recommendation
The next basket recommendation (NBR) task uses information from previous sessions. It is defined as recommending a group of items to a user based on their shopping history, where the history is a time-ordered sequence of baskets that they have purchased in the past [li-2023-next]. Each basket is a set of items with no particular order. This formulation fits the grocery shopping setting well, where a user’s purchase history occurs naturally in the form of such baskets. Variations of the NBR task where the order of items in a basket may be relevant, can, e.g., be found in music (playlist recommendation), in travel (recommending holiday packages), and in research and education (recommending reading lists). Two main characteristics of the grocery shopping scenario make the NBR in this domain distinct from other retail domains: (i) users shop for grocery items repeatedly and on a regular basis, and (ii) grocery items have a short life time and are repurchased frequently by the same user [liu-2019-characterizing].
In the grocery shopping domain, it has been found to be useful to distinguish between repeat items, i.e., items that a user has consumed before, and explore items, i.e., items that a user has not consumed before [ariannezhad-2022-recanet, li-2023-next]. In particular, for repeat items, the set of candidate items that needs to be considered for an individual user is usually in the low hundreds [ariannezhad-2021-understanding] as opposed to the full item catalog that needs to be considered for explore items, whose size may exceed 50,000 items in grocery shopping [ariannezhad-2020-demand, sprangers-2023-parameter]. This fact makes the task of retrieving repeat items to be included in the next basket to be recommended to a user considerably easier than the task of retrieving explore items. Frequency-based, nearest neighbor-based, and deep learning-based methods have all been used for the NBR task, and for the candidate retrieval phase in particular [li-2023-next]. As a rule of thumb, being biased towards the easier repetition task is an important strategy that helps to boost the overall NBR performance. Deep learning-based methods do not effectively exploit the repetition behavior. Indeed, they achieve a relatively good exploration performance, but they are not able to outperform the simple frequency-based baselines in several cases. Some recent state-of-the-art NBR methods are skewed towards the repetition task and outperform frequency-based baselines. However, the improvements they achieve are limited, especially considering the complexity and computational costs, e.g., for the training process [yu-2020-predicting] and for hyper-parameter search [faggioli-2020-recency, hu-2020-modeling].
@@ 6.3 Candidate ranking models
Given generic feature vectors as input, work on candidate ranking strategies has mainly focused on modeling interactions between features. According to the types of interaction function they adopt, existing methods can be divided into linear models, polynomial models, and neural network models.
@@ 6.3.1 Linear models
Early studies on candidate ranking usually apply linear models, such as logistic regression (LR) [kleinbaum2002logistic, hosmer2013applied] and naive Bayesian methods [hastie2009elements]. In contrast to other complicated models, linear models are straightforward, efficient, and explainable. Although linear models may not perform as well as deep neural networks, they indeed lay the foundation for recent advances in e-commerce recommendation [peng2002introduction, kiseleva2016beyond, bernardi2015continuous]. In e-commerce recommendation, logistic regression (LR) is one of the most popular methods that formulate the task as a classification task to rank through predicting the probability of an item to be interacted. LR first collects features of users (e.g., age and gender) and items (e.g., price and categories) into a number of feature vectors, and then apply a linear combination function to map the feature vector into the final predicted score. Similarly, kiseleva2016beyond employed a naive Bayesian ranking strategy in e-commerce recommendation by considering contextual user profiling.
@@ 6.3.2 Polynomial models
Performances of linear models are limited because of high space complexity and inability of high-level feature modeling. To address these two problems, factorization machines (FM) have been proposed [FM]. Factorization machines factorize parameters $w_{i,j}$ into an inner product of two latent vectors, i.e., $w_{i,j}\equiv\left<v_{i},v_{j}\right>$, where $v_{i}$ denotes a latent vector of the $i$-th feature. With different types of knowledge, the feature interactions across multiple fields should have different weights in recommendation. To tackle this challenge, field-aware factorization machines (FFM) [FFM] was proposed to capture field-aware weights and distribute a single latent vector to multiple fields.
A drawback of FM or FFM is that they just capture 2-order feature interactions but neglect higher order interactions, which are widely observed in e-commerce scenarios. As described in Section 4.1.1he2014practical proposed a hybrid model by combining GBDT and logistic regression for click-behavior modeling. The hybrid model is able to leverage the boosted decision trees (i.e., GBDT) to conduct feature interactions into logistic regression for e-commerce recommendation. In this hybrid model, GBDT adaptively conducts feature selection and higher-order feature interactions.
@@ 6.3.3 Neural network models
Deep neural networks have been leveraged in e-commerce recommendation because of their powerful expression ability of capturing complicated feature interactions. Up to now, research on neural network models can be categorized into the following three research directions:
  • The first direction aims at developing feature interaction modules based on neural networks, e.g., adding more neural network layers or combining the superiority of different neural networks.
  • The second direction aims at enhancing the expression by leveraging deep neural networks using FMs.
  • The third direction aims at leveraging the attention mechanism in capturing diverge and dynamic contributions of feature interactions.
In the upcoming sections, we will detail the recent advances along these research directions.
@@
Neural feature interactions. DeepCross proposed Deep Crossing model, which can be considered as the first end-to-end deep learning framework for recommender systems. Deep Crossing enjoys the merits of deep learning in coping with various features and capturing complex feature interactions, consisting of the following four components: the embedding layer, stacking layer, multiple residual units layer, and scoring layer. The goal of the embedding layer is to transform per individual sparse features into dense vectors in latent space via neural networks, whereas the stacking layer concatenates different embedding features from the embedding layer and generates a new vector for all features. The scoring layer servers as an output layer with logistic regression to generate the final predicted score.
@@
Collaborative filtering can be reconsidered from the perspective of deep learning. Traditional collaborative filtering methods employ the inner product of a user’s latent vector and an item’s latent vector for rating prediction. However, the simple model structure of MF makes it insufficient to fit the objective, which limits the expressiveness of the model. Motivated by this, neural collaborative filtering (NCF) [He2017NCF] was proposed to replace the inner product operation with a neural network.
@@
By learning frequent co-occurrences of features, deep neural networks may have poor memorization, i.e., these models easily over-generalize and recommend less relevant items when user-item interactions are sparse. To address this problem, cheng2016wide proposed the Wide&Deep model for recommendation. The detailed model architecture of Wide&Deep is presented in Section 4.1.1. Wide&Deep can keep a balance between memorization and generalization: its wide component can effectively memorize sparse feature interactions, while the deep neural networks can generalize the previously unseen feature interactions through low-dimensional embeddings.
@@
There are lots of follow-up studies that improve either the wide component or the deep component in the Wide&Deep model. Deep&Cross Network (DCN) [wang2017deep] replaces the wide component with a well-designed Cross network. DCN applies an explicit feature crossing mechanism with multiple cross layers. Recent studies seek to apply automating machine learning (AutoML) to model the selection of feature interactions [Su2021Detecting, Ze2021A, ZhaO2021AutoLoss]. SIF [Yao2020Efficient] utilizes one-shot architecture search methods to search proper interaction functions (e.g., inner product, plus/minor, max/min pooling, outer product, and concatenation) for collaborative filtering models. AutoFIS [Liu2020AutoFIS] continuously searches the effective feature interactions by incorporating architecture parameters to identify important feature interactions. AutoGroup [Liu2020AutoGroup] groups useful features into sets using AutoML and then generates interactions from each set.
@@
Endowing factorization machines with neural networks. Many models have been proposed to integrate factorization machines with deep neural networks to make full use of their advantages in feature combination. Neural factorization machines [he2017neural] enhances FMs by modeling higher-order and non-linear feature interactions. NFM introduced the bi-interaction pooling operation in neural network modeling, and presented a new neural network perspective for FMs. Given the embedding set of all features $\mathcal{V}_{x}$, the bi-interaction layer in NFM converts $\mathcal{V}_{x}$ to one vector:
@@
where $\odot$ denotes an element-wise product of two vectors. The output of bi-interaction pooling is a $d$-dimension vector that encodes the second-order interactions between features in the embedding space. Through stacking non-linear layers above the bi-interaction layer, NFM can effectively model higher-order and nonlinear feature interactions. In contrast to traditional deep learning methods that simply concatenate or average embedding vectors in the low level, the use of bi-interaction pooling encodes more informative feature interactions. Fig. 6.5 illustrates the architecture of NFM.
@@
As demonstrated in Section 4.1, DeepFM [guo2017deepfm] aims at learning both low and high-order feature interactions, and consists of two components: the FM component and the deep component. Compared with Wide&Deep, DeepFM replaces its wide component with FM to remedy its shortcoming in automatic feature combination. Another difference is that DeepFM shares the feature embedding between the FM and deep component. Besides NFM and DeepFM, many other neural networks also have been proposed based on FMs: FNN [zhang2016deep] directly stacks FMs with neural networks; PNN [qu2016product] models both bit-wise interactions and vector-wise feature interactions; and xDeepFM [lian2018xdeepfm] extends deepFM with explicit high-order feature interactions.
Attention mechanisms. Attention mechanisms have been applied in recommender systems and achieved a great success [AFM, LiRCRLM17, zhang2019deep]. As far as we know, AFM [AFM] is an early attempt to introduce the attention mechanism to recommendation. It can be regarded as an extension of NFM. The sum pooling operation in NFM treats all pairwise feature interactions equally, which may produce suboptimal results. To address this problem, AFM uses the attention mechanism on feature interactions by performing a weighted sum on the interacted vectors. The output of the attention-based pooling layer is projected into the prediction score. In session-based or sequential recommendation scenarios, LiRCRLM17 proposed an encoder-decoder model, namely neural attentive recommendation machine (NARM), to emphasize a user’s main purpose in the current session. The authors adopted a hybrid encoder structure with a global component for modeling long-term purposes and a local component for modeling short-term purposes. Based on the combined session representation, a bi-linear matching scheme was then applied to compute the recommendation scores for each candidate item. Following NARM, ren2018repeatnet proposed the RepeatNet model to deal with the phenomenon of repeat consumption behaviors. The authors incorporated a repeat-explore mechanism into neural networks, which can select items from a user’s history and suggests them at the appropriate moment. In standard embedding paradigms, user features are compressed into fixed-length representation vectors. However, the fixed-length vectors restrict the capture of the diverse interests of the user from historical behaviors. [zhou2018deep] proposed the deep interest network (DIN) to tackle this challenge by designing a local attention unit. The local attention unit in DIN adaptively learns the representation of user interests from historical behaviors by taking account of the relevance of historical behaviors. Yuan2021Looking unified existing CTR prediction models using a discrete choice model based on the self-attention mechanism. The authors regarded feature interaction as the individual’s comprehensive measurement of the influence of different factors on the decision-making process.

### v3

@@ 6.2.2 Embedding-based methods
Embedding-based methods are another type of candidate retrieval methods. As shown in Fig. 6.3, this kind of method first maps users and items into a common embedding space, and then uses approximate the K-Nearest Neighbor (KNN) algorithm [andoni2006near] to search for candidate items with the smallest embedding distance to the target user. The key of these methods lies in learning high-quality embeddings for each user and item. In what follows, we introduce techniques for learning embeddings.
Matrix factorization. Matrix factorization (MF) is a classic embedding-based method. The basic assumption behind MF is that the user-item interaction matrix has a low-rank structure. MF delineates each user and item as an embedding vector and then predicts the preference between each user-item pair as the inner product of their embedding vectors. Let $p_{u}\in\mathcal{R}^{k}$, $q_{i}\in\mathcal{R}^{k}$ denote the embedding vector of user $u$ and item $i$, respectively. MF makes a prediction for the user-item pair $(u,i)$ as follows:
@@
Compared with heuristic-based methods, MF is a more generic method that can adaptively learn user preferences from their history behaviors and require no manually crafted heuristic design. The use of MF has brought a revolution, pushing research attention from previous heuristic-based methods towards embedding-based methods.
Information-enhanced embedding models. Despite the prevalence of MF, it is still insufficient to yield accurate embeddings. The reason is that MF directly projects user/item IDs to an embedding space, making MF reliant on the behavioral signal from the objective function. Hence, MF models do not perform well for inactive users or items with very few interactions. To deal with this problem, several methods have been proposed to enrich the representation with supplementary information. We can divide them into three groups: (i) neighborhood-enriched embedding methods (ii) feature-enriched embedding methods and (iii) graph-enriched embedding methods. We discuss each of these groups in detail as follows:
(i) Neighborhood-enriched embedding. Beyond the user ID, bell2007lessons, koren2008svd++, kabbur2013fism enrich a user’s representation with their rating history. This kind of method generates user-item preference scores as follows:
@@
(ii) Feature-enriched embedding. Another way to improve MF is to use rich user and item features, e.g., a user’s age, gender, education, revenue, product tags, category, and price. These features are valuable in enriching the representations of users and items, which further boosts the recommendation performance. Fig. 6.4 summarizes the architecture of feature-enriched candidate retrieval methods as a two-tower structure [yi2019youtube-2tower, fan2019mobius, xu2016tag-dssm]. The left tower is the user tower that translates a user’s features into the user’s embedding representation; the right tower is the item tower generating the item representation. Let $x_{u}$ denote the features of user $u$, while $x_{i}$ denotes the features of item $i$. The function of the two towers can be depicted as follows:
@@
where $f_{U}(\cdot)$ and $f_{I}(\cdot)$ denote the translation function with regard to the implementation of the two towers. Linear models and neural networks (e.g., MLP [huang2013learning], LSTM [song2016TDDSSM], CNN [shen2014learning] and auto-encoders [wu2016collaborative, liang2018variational]) can be used as the translation function. Given user and item embeddings, the two-tower model makes a prediction for each user-item pair. Cosine similarity and inner product are usually adopted to measure the embedding distance, so we have:
@@
(iii) Graph-enriched embedding. A drawback of the aforementioned methods is that they regard each user or item as an “island” and fail to explicitly encode their relations into representations. These relations, e.g., item-item occurrences or user-item interactions, are important in revealing correlations between users and items, which provide valuable signals for recommendation. To address this problem, several studies have established specific graphs between users and items, and then conduct graph representation learning to generate user and item embeddings. One representative method is EGES [wang2018billion]. EGES constructs an item-item graph based on user behavior sequences, where two items are connected if they consecutively occur in one sequence. EGES then applies DeepWalk [perozzi2014deepwalk] on the item graph to generate item representations. Similarly, other related methods perform graph convolutional networks to enrich the representation learning of both users and items [wang2019NGCF, ying2018graph, LightGCN]. A knowledge graph with informative relations between items has been exploited to learn better embeddings [wang2019kgat, wang2020ckan].
@@ 6.2.3 Session-based recommendation
Early studies on sequential recommendation methods often use Markov chains by assuming that the user’s next action only depends on the previous one. As a representative method, the factorizing personalized Markov chains model (FPMC) extends MF by modeling the effects of sequential-consecutive actions [FPMC]. Since the previous action is a critical factor affecting the user’s next decision, FPMC achieves a significant gain over MF-based models. Following FPMC, he2016fusing use a higher-order Markov chain to capture the sequential dependence among non-consecutive behavior. It is hard to use Markov chain-based methods for capturing complicated and long-term dependencies in sequential data, and this limits their performance in e-commerce recommendation.
More recently, deep neural networks have been used in sequential recommendation due to their powerful expressive ability on capturing behavior dependences (discussed in Section 4). Generally speaking, neural networks for sequential recommendation can be divided into three types: (i) RNN-based methods that model sequential dependence with RNN (or improved versions such as LSTM and GRU) to capture both long-term and short-term dependencies [quadrana2017personalizing, jannach2017recurrent, hidasi2016parallel]; (ii) CNN-based methods that concatenate the embedding of the previous item in a sequence as to a matrix [tang2018personalized]; and (iii) attention-based methods that introduce attention mechanisms into sequential recommendation by considering various types of user behavior with different influences [kang2018self, sun2019bert4rec].
@@ 6.2.4 Next-basket recommendation
The next basket recommendation (NBR) task uses information from previous sessions. It is defined as recommending a group of items to a user based on their shopping history, where the history is a time-ordered sequence of baskets that they have purchased in the past [li-2023-next]. Each basket is a set of items with no particular order. This formulation fits the grocery shopping setting well, where a user’s purchase history occurs naturally in the form of such baskets. Variations of the NBR task where the order of items in a basket may be relevant, can, e.g., be found in music (playlist recommendation), in travel (recommending holiday packages), and in research and education (recommending reading lists). Two main characteristics of the grocery shopping scenario make the NBR task in this domain distinct from other retail domains: (i) users shop for grocery items repeatedly and on a regular basis, and (ii) grocery items have a short life time and are repurchased frequently by the same user [liu-2019-characterizing].
In the grocery shopping domain, it has been found to be useful to distinguish between repeat items, i.e., items that a user has consumed before, and explore items, i.e., items that a user has not consumed before [ariannezhad-2022-recanet, li-2023-next]. In particular, for repeat items, the set of candidate items that needs to be considered for an individual user is usually in the low hundreds [ariannezhad-2021-understanding] as opposed to the full item catalog that needs to be considered for explore items, whose size may exceed 50,000 items in grocery shopping [ariannezhad-2020-demand, sprangers-2023-parameter]. This fact makes the task of retrieving repeat items to be included in the next basket to be recommended to a user considerably easier than the task of retrieving explore items. Frequency-based, nearest neighbor-based, and deep learning-based methods have all been used for the NBR task, and for the candidate retrieval phase in particular [li-2023-next]. As a rule of thumb, being biased towards the easier repetition task is an important strategy that helps to boost the overall NBR performance. Deep learning-based methods do not effectively exploit the repetition behavior. Indeed, they achieve a relatively good exploration performance, but they are not able to outperform simple frequency-based baselines in several cases. Some recent state-of-the-art NBR methods are skewed towards the repetition task and outperform frequency-based baselines. However, the improvements they achieve are limited, especially considering the complexity and computational costs, e.g., for the training process [yu-2020-predicting] and for hyper-parameter search [DBLP:conf/um/FaggioliPA20, hu-2020-modeling].
@@ 6.3 Candidate ranking models
Given generic feature vectors as input, work on candidate ranking strategies has mainly focused on modeling interactions between features. According to the types of interaction function they adopt, existing methods can be divided into linear, polynomial, and neural network models.
@@ 6.3.1 Linear models
Early studies on candidate ranking usually apply linear models, such as logistic regression (LR) [kleinbaum2002logistic, hosmer2013applied] and naive Bayesian methods [hastie2009elements]. In contrast to other complicated models, linear models are straightforward, efficient, and explainable. Although linear models may not perform as well as deep neural networks, they indeed lay the foundation for recent advances in e-commerce recommendation [peng2002introduction, kiseleva2016beyond, bernardi2015continuous]. In e-commerce recommendation, logistic regression (LR) is one of the most popular methods that formulate the task as a classification task to rank through predicting the probability of an item to be interacted. LR first collects features of users (e.g., age and gender) and items (e.g., price and categories) into a number of feature vectors, and then applies a linear combination function to map the feature vector into the final predicted score. Similarly, kiseleva2016beyond employ a naive Bayesian ranking strategy in e-commerce recommendation by considering contextual user profiling.
@@ 6.3.2 Polynomial models
The performance of linear models is limited because of high space complexity and the inability of high-level feature modeling. To address these two problems, factorization machines (FM) have been proposed [FM]. Factorization machines factorize parameters $w_{i,j}$ into an inner product of two latent vectors, i.e., $w_{i,j}\equiv\left<v_{i},v_{j}\right>$, where $v_{i}$ denotes a latent vector of the $i$-th feature. With different types of knowledge, the feature interactions across multiple fields should have different weights in recommendation. To tackle this challenge, field-aware factorization machines (FFM) [FFM] have been proposed to capture field-aware weights and distribute a single latent vector to multiple fields.
A drawback of FM and FFM is that they only capture second-order feature interactions but neglect higher order interactions, which are widely observed in e-commerce scenarios. As described in Section 4.1.1he2014practical proposed a hybrid model by combining GBDT and logistic regression for click-behavior modeling. The hybrid model is able to use boosted decision trees (i.e., GBDT) to conduct feature interactions into logistic regression for e-commerce recommendation. In this hybrid model, GBDT adaptively conducts feature selection and higher-order feature interactions.
@@ 6.3.3 Neural network models
Deep neural networks have been used in e-commerce recommendation because of their powerful expression ability of capturing complicated feature interactions. Up to now, research on neural network models can be categorized into the following three research directions: (i) The first direction aims at developing feature interaction modules based on neural networks, e.g., adding more neural network layers or combining the superiority of different neural networks. (ii) The second direction aims at enhancing the expression by using deep neural networks using FMs. (iii) The third direction aims at using the attention mechanism in capturing diverge and dynamic contributions of feature interactions. Below, we will detail the recent advances along these research directions.
@@
Neural feature interactions. DeepCross propose the deep crossing model, which can be considered as the first end-to-end deep learning framework for recommender systems. Deep crossing enjoys the merits of deep learning in coping with various features and capturing complex feature interactions. It consists of the four components: an embedding layer, a stacking layer, multiple residual units layer, and a scoring layer. The goal of the embedding layer is to transform per individual sparse features into dense vectors in latent space via neural networks. The stacking layer concatenates different embedding features from the embedding layer and generates a new vector for all features. The scoring layer servers as an output layer with logistic regression to generate the final predicted score.
@@
Collaborative filtering can be reconsidered from the perspective of deep learning. Traditional collaborative filtering methods employ the inner product of a user’s latent vector and an item’s latent vector for rating prediction. Neural collaborative filtering (NCF) [He2017NCF] has been proposed to replace the inner product operation with a neural network.
@@
By learning frequent co-occurrences of features, deep neural networks may have poor memorization, i.e., these models easily over-generalize and recommend less relevant items when user-item interactions are sparse. To address this problem, cheng2016wide introduce the Wide&Deep model for recommendation. The detailed model architecture of Wide&Deep has already been presented in Section 4.1.1. Wide&Deep maintains a balance between memorization and generalization: its wide component can effectively memorize sparse feature interactions, while the deep neural networks can generalize the previously unseen feature interactions through low-dimensional embeddings.
@@
Follow-up studies improve either the wide component or the deep component in the Wide&Deep model. The Deep&Cross Network (DCN) [wang2017deep] replaces the wide component with a well-designed cross network. DCN applies an explicit feature crossing mechanism with multiple cross layers. Later studies seek to apply automating machine learning (AutoML) to model the selection of feature interactions [Su2021Detecting, Ze2021A, ZhaO2021AutoLoss]. SIF [Yao2020Efficient] uses one-shot architecture search methods to search proper interaction functions (e.g., inner product, plus/minor, max/min pooling, outer product, and concatenation) for collaborative filtering models. AutoFIS [Liu2020AutoFIS] continuously searches effective feature interactions by incorporating architecture parameters to identify important feature interactions. AutoGroup [Liu2020AutoGroup] groups useful features into sets using AutoML and then generates interactions from each set.
@@
Endowing factorization machines with neural networks. Many models have been proposed to integrate factorization machines with deep neural networks to make full use of their advantages in feature combination. Neural factorization machines [NFMs; he2017neural] enhance FMs by modeling higher-order and non-linear feature interactions. NFM introduces a bi-interaction pooling operation in neural network modeling, and presents a new neural network perspective for FMs. Given the embedding set of all features $\mathcal{V}_{x}$, the bi-interaction layer in NFM converts $\mathcal{V}_{x}$ to one vector:
@@
where $\odot$ denotes an element-wise product of two vectors. The output of bi-interaction pooling is a $d$-dimensional vector that encodes the second-order interactions between features in the embedding space. By stacking non-linear layers above the bi-interaction layer, NFM can effectively model higher-order and non-linear feature interactions. In contrast to traditional deep learning methods that simply concatenate or average embedding vectors in the low level, the use of bi-interaction pooling encodes more informative feature interactions. Fig. 6.5 illustrates the architecture of NFM.
@@
As demonstrated in Section 4.1, DeepFM [guo2017deepfm] aims to learn both low and high-order feature interactions, and consists of two components: the FM component and the deep component. Compared with Wide&Deep, DeepFM replaces its wide component with FM to remedy its shortcoming in automatic feature combination. Another difference is that DeepFM shares the feature embedding between the FM and deep component. Besides NFM and DeepFM, many other neural networks have been proposed based on FMs: FNN [zhang2016deep] directly stacks FMs with neural networks; PNN [qu2016product] models both bit-wise interactions and vector-wise feature interactions; and xDeepFM [lian2018xdeepfm] extends deepFM with explicit high-order feature interactions.
Attention mechanisms. Attention mechanisms have been applied in recommender systems and achieved great success [AFM, LiRCRLM17, zhang2019deep]. AFM [AFM] is an early attempt to introduce an attention mechanism to recommendation. It can be regarded as an extension of NFM. The sum pooling operation in NFM treats all pairwise feature interactions equally, which may produce suboptimal results. To address this problem, AFM uses the attention mechanism on feature interactions by performing a weighted sum on the interacted vectors. The output of the attention-based pooling layer is projected into the prediction score. In session-based or sequential recommendation scenarios, LiRCRLM17 propose an encoder-decoder model, neural attentive recommendation machine (NARM), to emphasize a user’s main purpose in the current session. The authors adopt a hybrid encoder structure with a global component for modeling long-term purposes and a local component for modeling short-term purposes. Based on the combined session representation, a bi-linear matching scheme is then applied to compute the recommendation scores for each candidate item. Following NARM, ren2018repeatnet put forward the RepeatNet model to deal with the phenomenon of repeat consumption behavior. The authors incorporate a repeat-explore mechanism into neural networks, which can select items from a user’s history and suggests them at the appropriate moment. In standard embedding paradigms, user features are compressed into fixed-length representation vectors. However, fixed-length vectors limit capturing the diverse interests of a user from historical behavior. [zhou2018deep] introduce the deep interest network (DIN) to tackle this challenge by designing a local attention unit. The local attention unit in DIN adaptively learns the representation of user interests from historical behavior by taking account of the relevance of historical behavior. Yuan2021Looking unify CTR prediction models using a discrete choice model based on the self-attention mechanism. The authors regard feature interaction as the individual’s comprehensive measurement of the influence of different factors on the decision-making process.